OpenStudy (anonymous):

Solve the system algebraically. Check your work. 2x + 5y = 10 2x + 3y = 6

OpenStudy (anonymous):

\[3y-5y=6-10\]

OpenStudy (anonymous):

\[-2y=-4\]

OpenStudy (anonymous):

\[y=2\]

OpenStudy (anonymous):

\[2x+5(2)=10\]

OpenStudy (anonymous):

\[2x=0 \]

OpenStudy (anonymous):

\[x=0\]